OpenStudy (anonymous):

Simplify. square root 75 minus 4 times the square root 8 minus 3 times the square root of 32 negative 15 square root 5 negative 15 square root 6 5 square root 3 minus 20 square root 2 5 square root 3 plus 4 square root 2

OpenStudy (anonymous):

square root 75 \[ \sqrt{75} \] square root 75 minus 4 times the square root 8 \[ \sqrt{75}-4\sqrt{8} \] square root 75 minus 4 times the square root 8 minus 3 times the square root of 32\[ \sqrt{75}-4\sqrt{8}-3\sqrt{32} \] Okay so far is this easy to understand?

OpenStudy (anonymous):

some what

OpenStudy (anonymous):

Now we factor: \[ 75 = 3\times 5\times 5 \] Can you factor \(8\) and \(32\)?

OpenStudy (anonymous):

Notice how:\[ \sqrt{75} =\sqrt{3\times 5^2}=5\sqrt{3} \]

OpenStudy (anonymous):

We want to do this with \(8\) and \(32\) if possible.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

@wio how do you want me to do this?

OpenStudy (anonymous):

First factor the numbers \(8\) and \(32\) into prime numbers.

OpenStudy (anonymous):

ok i dont undersyand the factorin part

OpenStudy (anonymous):

Basically, what two numbers do you multiply to get \(8\)?

OpenStudy (anonymous):

2&4

OpenStudy (anonymous):

Ok, to do 8,